# A metrological framework for uncertainty evaluation in machine learning classification models

### v2

@@ 4.1.1 Background
  • Hydrological and climate modelling [Bohn2019, Wang2023], where LC maps are used for setting up land surface conditions;
  • National greenhouse gas accounting [IPCC2006], where LC maps along with additional data are used for estimation of emissions associated with anthropogenic land use;
  • Biodiversity monitoring [Skidmore2021], where LC maps provide information on habitats and their fragmentation, etc.
@@
LC mapping is a multi-class classification problem, which currently is widely approached by ML. The most popular ML classification models used for LC mapping are tree-based models (such as random forests) [Kilcoyne2022, Marston2022, Friedl2022, Venter2021], and convolutional neural networks (CNNs) [Brown2022, Karra2021]. There are many LC classification schemes. Some classification schemes include only a few, very broad, classes, e.g., tree cover, grassland, cropland, etc. [Zanaga2022], whereas other classification schemes provide finer details, e.g. they distinguish broad-leaved and coniferous forest [EEA2020]. Some classification schemes have a nested structure where broad classes are further split into narrower classes [Friedl2022]. It is important to note that in the present case study we approach land cover classification in similar fashion to how it is commonly approached in the literature – as a classification problem with a univariate output, i.e., each pixel considered separately. An important extension would be to perform classification with a multivariate output. In this case, classification would produce a collection of output variables (a collection of LC classes) corresponding to multiple pixels as opposed to each single pixel. Such an extension would make sense given how strongly correlated EO data typically is in the spatial domain.
@@ 4.1.2 Data
As previously mentioned in Section 1.1, a supervised ML classification model is trained on a data set consisting of pairs of several input variables and a categorical output. In our LC classification study, the input variables represented surface reflectance measurements performed in ten spectral bands of Sentinel-2 satellite over a single pixel. The output variable represented LC class in that pixel. Information on LC classes (i.e., LC labels) is typically obtained either manually (by visual interpretation of high/very high resolution satellite imagery or by ground surveys), automatically (e.g., based on pre-existing LC maps), or hybrid-like (following a combination of the two methods) [Moraes2024]. In our study, LC labels were obtained by an automatic collection procedure following a modified approach of UK Centre for Ecology and Hydrology (UKCEH), used for the collection of training data for the production of UK annual LC maps [Marston2022]. This method relies on the assumption that LC change is gradual. It uses UKCEH LC maps produced for three previous years to find pixels with stable LC classes (i.e. unchanged LC classes throughout these years). Further filtering was performed to identify such unchanged pixels where the RF classifier used by the UKCEH to produce their LC maps yielded a probability greater than 95 %. In other words, we only retained pixels, where the LC label was assigned with the highest confidence. For the present case study, LC labels were collected over a $20\,\text{km}\times 20\,\text{km}$ area of interest (AOI) in Scotland. The collection process yielded 189,142 unchanged pixels (i.e., 4.7 % of the AOI) corresponding to forest, cropland, grassland and settlements classes.
The input variables (i.e., surface reflectance measurements) for the 189,142 unchanged pixels were retrieved from Sentinel-2 images collected over the AOI on the 1st of June, 2020 and 2021. Although space-borne surface reflectance measurements have many uncertainty sources associated with them, in the present study we choose to focus on one uncertainty source. This uncertainty source is related to the imperfect knowledge of aerosol optical depth (AOD) used in the atmospheric correction of the satellite imagery. It represents just one (although ordinarily dominant) contribution to the uncertainty sources (1) and (5) described in Section 1.2 and illustrated in Figure 1. The AOD uncertainties were introduced into the atmospheric correction process by Monte Carlo (MC) sampling (we used 25 MC samples). This process yielded 25 realizations of surface reflectance measurements for each of the 189,142 unchanged pixels. See [bilson2025uncertaintyawarebayesianmachinelearning] for more details on the data set.
@@ 4.1.3 Propagation of uncertainties within Bayesian classification models
As illustrated in Figure 1, ML classification models come with multiple sources of uncertainty. Most of these sources can be described in terms of probability distributions for particular parametric classification models (such as Naïve Bayes, logistic regression, discriminant analysis [murphy2012machine] and neural networks [jospin_hands-bayesian_2022]) that naturally lend themselves to a Bayesian treatment. In this section we describe a framework that can be followed to incorporate these sources of uncertainty for Bayesian generative ML classifiers, which explicitly model the input data uncertainties.
@@
  • ML models also come with model hyperparameters, which describe the choice of model (3), such as model architecture or optimisation method. Various hyperparameter optimisation tuning methods [franceschi2024hyperparameteroptimizationmachinelearning] can be used to determine the optimal model choice. Bayesian optimisation is one popular approach [NIPS2012_05311655], and for such an approach, quantification of the uncertainties in choice of model has been attempted [pmlr-v151-tuo22a].
@@
    $\mathcal{L}(\mathcal{D}|\theta)=\prod_{i=1}^{N}p(\mathbf{x}_{i},y_{i}|\theta)=\prod_{i=1}^{N}p(y_{i}|\mathbf{x}_{i},\theta)p(\mathbf{x}_{i}).$ (4.2)
@@
    $p(\theta|\mathcal{D})=\dfrac{\mathcal{L}(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})}.$ (4.3)
    The Bayesian approach introduces an additional uncertainty $p(\theta)$, which is the prior distribution on the model parameters. This can be determined from expert knowledge, or one can use conjugate priors [murphy2012machine] if known for analytic determination of the posterior. $p(\mathcal{D})$ is known as the evidence, and is typically intractable for more complex classification models such as CNNs. In such cases one must resort to sampling techniques such as Markov Chain Monte Carlo (MCMC) [brooks_handbook_2011], approximation techniques such as variational inference [murphy2012machine] or MC dropout [kendall2017uncertainties] to estimate the posterior distribution.
@@ 4.1.4 Numerical examples
We apply the Bayesian framework described in Section 4.1.3 to the datasets described in Section 4.1.2, using the Bayesian quadratic discriminant analysis model presented in [bilson2025uncertaintyawarebayesianmachinelearning]. Specifically we compute the posterior predictive probabilities on test sets ($90\,\%$ of available pixels) from years 2020 and 2021, with training set taken from the remaining $10\,\%$ of pixels from year 2020.
@@
For evaluation of the performance of the output class probabilities from the classification models, we use two expected values of proper scoring rules [gneiting2007strictly], namely the multi-class cross-entropy loss (XE) and the Brier score (BS). Proper scoring rules are a principled way to measure the quality of the predictive probabilities produced by ML classification models [ferrer_analysis_2023].
@@
The resulting metric values for test year 2020 were: classification loss = 0.0122; EXE = 0.0789; EBS = 0.0312. For test year 2021: classification loss = 0.0572; EXE = 0.5673; EBS = 0.1514. As is common with ML classification tasks, we see a degradation in performance when the model is tested on data outside the training domain (in this case a different year). We also observe from Figure 4 that the class ‘settlements’ have the highest false positive rate. This is in large part due to the low abundance of settlement pixels within the AOI, and so the ML classifier will tend to give other land cover classes with larger abundances a higher predictive probability.
@@
Figure 4: Confusion matrices using Bayesian QDA. Final column corresponds to the true/false positive rates for each class. Final row corresponds to true/false negative rates.
In addition to the classification results, we also compute the expressions of uncertainty for nominal properties described in Section 3.2 using the posterior PMFs obtained within the Bayesian framework of [bilson2025uncertaintyawarebayesianmachinelearning]. Figure 5 shows box and whisker plots [du_toit_graphical_1986] in the chosen uncertainty statistics, capturing the median, interquartile range (IQR), as well the limit of non-outliers (given by $1.5\times\text{IQR}$ away from the top or bottom box). Due to the high accuracy of the LC classifiers, all statistics had values very close to zero, so the median cannot be visualised [^2] Except for $H$ in test year 2021.. This also makes the distributions in the statistics highly skewed.
(a) Test Year: 2020
(b) Test Year: 2021
Figure 5: Box and whisker plots of uncertainty statistics for LC classification using Bayesian QDA.
Note the large difference in scales in Figure 5 from years 2020 to 2021. This is due to the large classification performance degradation observed in the classification metrics from years 2020 to 2021 (see Figure 4), causing the resulting predictive PMFs to be more uncertain. However, the uncertainty statistics still show high certainty in comparison to a uniform distribution. We also see a clear pattern for both years, with the information entropy $H$ having the largest IQR, and UVR having the smallest. Thus the entropy is the most sensitive to changes in the PMF out of all the uncertainty statistics considered. Comparing with Table 2, the statistics with the lowest IQR correspond to those that explicitly use the modal probability.
@@ 4.2.1 Background
Atrial Fibrillation (AF) is a heart condition that causes an irregular and often fast heart rate. The global prevalence of AF was estimated to be 59 million individuals in 2019. Patients with AF are at increased risk of stroke, heart failure and dementia, and it is considered to be a 21st century cardiovascular disease epidemic [linz2024atrial].
Digital health solutions are often used for the timely detection of AF, and diagnosis of AF is typically based upon an ECG. Wearable or smartphone-based photoplethysmography (PPG) technology has more recently emerged as having clear potential to make continuous monitoring of AF more widely available. PPG measurements use a light source and a detector on the surface of the skin to measure changes in the optical properties of the skin’s microvascular bed caused by relative changes in blood volume [bench2024towards].
Detection of AF from a time series segment of PPG measurements can be posed as a binary classification problem with two categories: positive (AF is present) and negative (AF is not present). Promising results for this classification task have already been reported. For example, a 122-patient study in [Wouters2024ComparativeEO] reported a specificity (Type I error) of 98.9 % and a sensitivity (Type II error) of 100 % for Fibricheck’s PPG-based AF detection smartphone app [fibricheck].
@@ 4.2.2 Data
We make use of a labelled PPG dataset called DeepBeat introduced in [torres2020multi]. A custom subset of the dataset was selected to ensure a balanced label distribution and no overlap of patients between the training and test datasets. A total of 134 patients were selected, giving a derived dataset consisting of over 100,000 signal segments, with over 50,000 instances of both AF and non-AF. Each segment has a duration of 25 seconds and a sampling rate of 32 Hz. See [torres2020multi] for further details on the primary dataset, and see [bench2024towards] for further details on the derived dataset.
@@ 4.2.3 Probabilistic ML model
There are generally two approaches to the AF classification task [charlton2022wearable]: (i) a rule-based classifier which uses features relating to pulse wave shape or inter-beat intervals; or (ii) an ML-based classifier which uses a training set of PPG data labelled by a clinical expert to optimise a generic empirical classification model (for example a neural network). This case study focuses on the latter approach.
@@
Recall that the dataset described in [torres2020multi] consists of signal segments from 134 patients. In [bench2024towards], data from 110 of the patients was used to train the model, which allows for data from the remaining 24 patients to be used as examples of the use of the trained model for probabilistic AF classification.
@@
We give a brief overview of the ML model and its optimisation, and refer the interested reader to [bench2024towards] for further details. We use a CNN model which is a close variant of xresnet1d50 reported in [strodthoff2020deep], which is itself a 1D variant of $\mathtt{xresnet}$ [he2019bag]. A CNN model consists of a sequence of linear transformations (convolutions) interspersed with simple non-linear operations, and it outputs prediction probabilities for each of the classes. Such a model is parametrised by a very large number of weights and biases defining each of the convolutions, and training the model involves optimising these parameters with respect to some loss function.
@@
The probabilities returned by a CNN capture contributions to the prediction uncertainty arising from some of the sources of uncertainty mentioned in Section 1, namely the class assignment uncertainty (2), the training data uncertainty (1) and the input test data uncertainty (5). These sources of uncertainty are those which are properties of the classification task itself or the measurement system used to obtain the observations. In the ML community, such sources of uncertainty are often referred to as aleatoric. Given observations, this is the contribution to the prediction uncertainty that cannot be reduced by choosing a better model or obtaining additional training samples. Aleatoric uncertainty is often distinguished in ML from epistemic uncertainty, which is the uncertainty concerning the choice (3) and optimisation (4) of the model.
@@
In order to capture in addition contributions to the prediction uncertainty due to model parameter uncertainty (4), an approach known as Monte Carlo Dropout [gal_dropout_2016, kendall2017uncertainties] is used. The approach involves enforcing convolution weights to be zero at random with some fixed probability when both training and evaluating the model. A probability distribution is finally obtained which combines the aleatoric and epistemic contributions to the prediction uncertainty [kendall2017uncertainties].
@@
We note two sources of uncertainty which are inadequately captured by the Monte Carlo Dropout approach. Firstly, as is usually the case for uncertainty evaluation of ML models, uncertainty concerning the choice of model is ignored. It follows therefore that reliable uncertainty evaluation assumes the selection of an adequate type of model. Secondly, only uncertainty of the output data is explicitly modelled in this approach, and uncertainty concerning input data is not explicitly modelled. This is in contrast to the Bayesian QDA approach in Section 4.1. Approaches which explicitly model uncertainty of the inputs to deep learning models have been considered [martin2023aleatoric], but the approach is computationally expensive and consequently not practically feasible for the large-scale AF detection task considered here.

### v3

@@ 4.1.1 Background
  • Hydrological and climate modelling [54, 55], where LC maps are used for setting up land surface conditions;
  • National greenhouse gas accounting [56], where LC maps along with additional data are used for estimation of emissions associated with anthropogenic land use;
  • Biodiversity monitoring [57], where LC maps provide information on habitats and their fragmentation, etc.
@@
LC mapping is a multi-class classification problem, which currently is widely approached by ML. The most popular ML classification models used for LC mapping are tree-based models (such as random forests) [58, 59, 60, 61], and convolutional neural networks (CNNs) [62, 63]. There are many LC classification schemes. Some classification schemes include only a few, very broad, classes, e.g., tree cover, grassland, cropland, etc. [64], whereas other classification schemes provide finer details, e.g. they distinguish broad-leaved and coniferous forest [65]. Some classification schemes have a nested structure where broad classes are further split into narrower classes [60]. It is important to note that in the present case study we approach land cover classification in similar fashion to how it is commonly approached in the literature – as a classification problem with a univariate output, i.e., each pixel considered separately. An important extension would be to perform classification with a multivariate output. In this case, classification would produce a collection of output variables (a collection of LC classes) corresponding to multiple pixels as opposed to each single pixel. Such an extension would make sense given how strongly correlated EO data typically is in the spatial domain.
@@ 4.1.2 Data
As previously mentioned in Section 1.1, a supervised ML classification model is trained on a data set consisting of pairs of several input variables and a categorical output. In our LC classification study, the input variables represented surface reflectance measurements performed in ten spectral bands of Sentinel-2 satellite over a single pixel. The output variable represented LC class in that pixel. Information on LC classes (i.e., LC labels) is typically obtained either manually (by visual interpretation of high/very high resolution satellite imagery or by ground surveys), automatically (e.g., based on pre-existing LC maps), or hybrid-like (following a combination of the two methods) [66]. In our study, LC labels were obtained by an automatic collection procedure following a modified approach of UK Centre for Ecology and Hydrology (UKCEH), used for the collection of training data for the production of UK annual LC maps [59]. This method relies on the assumption that LC change is gradual. It uses UKCEH LC maps produced for three previous years to find pixels with stable LC classes (i.e. unchanged LC classes throughout these years). Further filtering was performed to identify such unchanged pixels where the RF classifier used by the UKCEH to produce their LC maps yielded a probability greater than 95 %. In other words, we only retained pixels, where the LC label was assigned with the highest confidence. For the present case study, LC labels were collected over a $20\,\text{km}\times 20\,\text{km}$ area of interest (AOI) in Scotland. The collection process yielded 189,142 unchanged pixels (i.e., 4.7 % of the AOI) corresponding to forest, cropland, grassland and settlements classes.
The input variables (i.e., surface reflectance measurements) for the 189,142 unchanged pixels were retrieved from Sentinel-2 images collected over the AOI on the 1st of June, 2020 and 2021. Although space-borne surface reflectance measurements have many uncertainty sources associated with them, in the present study we choose to focus on one uncertainty source. This uncertainty source is related to the imperfect knowledge of aerosol optical depth (AOD) used in the atmospheric correction of the satellite imagery. It represents just one (although ordinarily dominant) contribution to the uncertainty sources (1) and (5) described in Section 1.2 and illustrated in Figure 1. The AOD uncertainties were introduced into the atmospheric correction process by Monte Carlo (MC) sampling (we used 25 MC samples).The process yielded 25 realizations of surface reflectance measurements for each of the 189,142 unchanged pixels. The relatively low number of MC samples used in this study reflects its preliminary nature. The current implementation serves as a proof-of-concept, and we plan to increase both the number of MC samples and the range of uncertainty sources in future iterations of the framework. See [17] for more details on the data used in this study.
@@ 4.1.3 Propagation of uncertainties within Bayesian classification models
As illustrated in Figure 1, ML classification models come with multiple sources of uncertainty. Most of these sources can be described in terms of probability distributions for particular parametric classification models (such as Naïve Bayes, logistic regression, discriminant analysis [16] and neural networks [67]) that naturally lend themselves to a Bayesian treatment. In this section we summarise a framework that can be followed to incorporate these sources of uncertainty for Bayesian generative ML classifiers, which explicitly model the input data uncertainties. For specific details of the framework and its implementation on land cover data, with reference to various uncertainty sources, see [17].
@@
  • ML models also come with model hyperparameters, which describe the choice of model (3), such as model architecture or optimisation method. Various hyperparameter optimisation tuning methods [68] can be used to determine the optimal model choice. Bayesian optimisation is one popular approach [69], and for such an approach, quantification of the uncertainties in choice of model has been attempted [70].
@@
    $\mathcal{L}(\theta|\mathcal{D})=p(\mathcal{D}|\theta)=\prod_{i=1}^{N}p(\mathbf{x}_{i},y_{i}|\theta)=\prod_{i=1}^{N}p(y_{i}|\mathbf{x}_{i},\theta)p(\mathbf{x}_{i}).$ (4.2)
@@
    $p(\theta|\mathcal{D})=\dfrac{p(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})}.$ (4.3)
    The Bayesian approach introduces an additional uncertainty $p(\theta)$, which is the prior distribution on the model parameters. This can be determined from expert knowledge, or one can use conjugate priors [16] if known for analytic determination of the posterior. $p(\mathcal{D})$ is known as the evidence, and is typically intractable for more complex classification models such as CNNs. In such cases one must resort to sampling techniques such as Markov Chain Monte Carlo (MCMC) [71], approximation techniques such as variational inference [16] or MC dropout [19] to estimate the posterior distribution.
@@ 4.1.4 Numerical examples
We apply the Bayesian framework described in Section 4.1.3 to the datasets described in Section 4.1.2, using the Bayesian quadratic discriminant analysis (BQDA) model presented in [17]. Specifically we compute the posterior predictive probabilities on test sets ($90\,\%$ of available pixels) from years 2020 and 2021, with training set taken from the remaining $10\,\%$ of pixels from year 2020.
@@
For evaluation of the performance of the output class probabilities from the classification models, we use two expected values of proper scoring rules [72], namely the multi-class cross-entropy loss (XE) and the Brier score (BS). Proper scoring rules are a principled way to assess the quality of the predictive probabilities produced by ML classification models [73].
@@
4.1.5 Results
The resulting metric values for assessment of the performance of the classifiers are given in Table 3. As is common with ML classification tasks, we see a degradation in performance when the model is tested on data outside the training domain (in this case a different year). We also observe from Figure 4 that the class ‘Settlements’ have the highest false positive rate. This is in large part due to the low abundance of settlement pixels within the AOI, and so the ML classifier will tend to give other land cover classes with larger abundances a higher predictive probability.
Year Loss EXE EBS
2020 0.012 0.079 0.031
2021 0.057 0.567 0.151
Table 3: Performance metrics for LC classification using BQDA model
@@ 4.1.5 Results
Figure 4: Confusion matrices using BQDA. Final column corresponds to the true/false positive rates for each class. Final row corresponds to true/false negative rates.
In addition to the classification results, we also compute the expressions of uncertainty for nominal properties described in Section 3.2 using the posterior PMFs obtained within the Bayesian framework of [17]. Tables 4 and 5 show a summary of the behaviour in the chosen uncertainty statistics, for test years 2020 and 2021 respectively.
[…]
WVR $4.63$ $8.10$ $2.72$ $5.00$
UVR $3.71$ $6.48$ $2.18$ $4.00$
SDM $4.63$ $7.81$ $2.72$ $4.72$
$H$ $88.8$ $12.9$ $24.6$ $5.84$
$H^{*}$ $41.0$ $7.35$ $11.3$ $3.63$
IQV $9.26$ $12.4$ $5.45$ $6.73$
CNV $4.63$ $7.04$ $2.72$ $4.02$
Table 4: Summary of uncertainty statistics for LC classification. Test Year: 2020
[…]
WVR $7.44$ $3.03$ $3.16$ $9.94$
UVR $5.95$ $2.42$ $2.53$ $7.95$
SDM $7.44$ $2.91$ $3.16$ $9.34$
$H$ $51.9$ $4.33$ $13.2$ $11.1$
$H^{*}$ $24.0$ $2.56$ $6.15$ $7.17$
IQV $14.9$ $4.48$ $6.31$ $13.1$
CNV $7.44$ $2.59$ $3.16$ $7.97$
Table 5: Summary of uncertainty statistics for LC classification. Test Year: 2021
In both tables, we see that the median and mean values are very small. This is due to the high accuracy or confidence of the LC classifiers, leading to low uncertainty. In both tables, we also see a pattern that the median in orders of magnitude smaller than the mean, and the IQR is orders of magnitude smaller than the s.d. This is because the distributions of the uncertainty statistics are highly skewed to the left, as again most of the classifier predictions were highly confident, leading values of the uncertainty statistics to be close to zero.
Note the large difference in scales between Tables 4 and 5. This is due to the large classification performance degradation observed in the classification metrics from years 2020 to 2021 (see Figure 4), causing the resulting predictive PMFs to be more uncertain. However, the uncertainty statistics still show high certainty in comparison to a uniform distribution. We also see a clear pattern for both years, with the information entropy $H$ having the largest median and IQR, and UVR having the smallest. Since the median and IQR represent the behaviour of the statistics for small values, the entropy is the most sensitive to variations in confident PMFs out of all the uncertainty statistics considered. UVR is the least sensitive in this case. Alternatively, the IQV has the largest s.d., and largest mean for year 2021, indicating that this statistic is more sensitive to variations in uncertain PMFs, i.e. PMFs that are close to uniform. The transformed entropy $H^{*}$ is the least sensitive, or most robust against small changes in uncertain PMFs.
We observe that WVR, SDM, and CNV have equivalent median and IQR in Tables 4 and 5. This is because the median and IQR are describing PMFs with a large $\widehat{p}$ in comparison with the other classes. This reduces these expressions for the uncertainty statistics of (3.16)(3.25) to approximately (3.12), essentially acting as a binary classification PMF where all other classes are treated as one. Appendix B gives a derivation of this equivalence.
@@ 4.2.1 Background
[…]
Digital health solutions are often used for the timely detection of AF, and diagnosis of AF is typically based upon an ECG. Wearable or smartphone-based photoplethysmography (PPG) technology has more recently emerged as having clear potential to make continuous monitoring of AF more widely available. PPG measurements use a light source and a detector on the surface of the skin to measure changes in the optical properties of the skin’s microvascular bed caused by relative changes in blood volume [75].
Detection of AF from a time series segment of PPG measurements can be posed as a binary classification problem with two categories: positive (AF is present) and negative (AF is not present). Promising results for this classification task have already been reported. For example, a 122-patient study in [76] reported a specificity (Type I error) of 98.9 % and a sensitivity (Type II error) of 100 % for Fibricheck’s PPG-based AF detection smartphone app [77].
@@ 4.2.2 Data
We make use of a labelled PPG dataset called DeepBeat introduced in [7]. A custom subset of the dataset was selected to ensure a balanced label distribution and no overlap of patients between the training and test datasets. A total of 134 patients were selected, giving a derived dataset consisting of over 100,000 signal segments, with over 50,000 instances of both AF and non-AF. Each segment has a duration of 25 seconds and a sampling rate of 32 Hz. See [7] for further details on the primary dataset, and see [75] for further details on the derived dataset.
@@ 4.2.3 Probabilistic ML model
There are generally two approaches to the AF classification task [78]: (i) a rule-based classifier which uses features relating to pulse wave shape or inter-beat intervals; or (ii) an ML-based classifier which uses a training set of PPG data labelled by a clinical expert to optimise a generic empirical classification model (for example a neural network). This case study focuses on the latter approach.
@@
Recall that the dataset described in [7] consists of signal segments from 134 patients. In [75], data from 110 of the patients was used to train the model, which allows for data from the remaining 24 patients to be used as examples of the use of the trained model for probabilistic AF classification.
@@
We give a brief overview of the ML model and its optimisation, and refer the interested reader to [75] for further details. We use a CNN model which is a close variant of xresnet1d50 reported in [79], which is itself a 1D variant of $\mathtt{xresnet}$ [80]. A CNN model consists of a sequence of linear transformations (convolutions) interspersed with simple non-linear operations, and it outputs prediction probabilities for each of the classes. Such a model is parametrised by a very large number of weights and biases defining each of the convolutions, and training the model involves optimising these parameters with respect to some loss function.
@@
The probabilities returned by a CNN capture contributions to the prediction uncertainty arising from some of the sources of uncertainty mentioned in Section 1, namely the class assignment uncertainty (2), the training data uncertainty (1) and the input test data uncertainty (5). These sources of uncertainty are those which are properties of the classification task itself or the measurement system used to obtain the observations. In the ML community, such sources of uncertainty are often referred to as aleatoric. Given observations, this is the contribution to the prediction uncertainty that cannot be reduced by choosing a better model or obtaining additional training samples. Aleatoric uncertainty is often distinguished in ML from epistemic uncertainty, which is the uncertainty concerning the choice (3) and optimisation (4) of the model [15].
@@
In order to capture in addition contributions to the prediction uncertainty due to model parameter uncertainty (4), an approach known as Monte Carlo Dropout [18, 19] is used. The approach involves enforcing convolution weights to be zero at random with some fixed probability when both training and evaluating the model. A probability distribution is finally obtained which combines the aleatoric and epistemic contributions to the prediction uncertainty [19].
@@
We note two sources of uncertainty which are inadequately captured by the Monte Carlo Dropout approach. Firstly, as is usually the case for uncertainty evaluation of ML models, uncertainty concerning the choice of model is ignored. It follows therefore that reliable uncertainty evaluation assumes the selection of an adequate type of model. Secondly, only uncertainty of the output data is explicitly modelled in this approach, and uncertainty concerning input data is not explicitly modelled. This is in contrast to the Bayesian QDA approach in Section 4.1. Approaches which explicitly model uncertainty of the inputs to deep learning models have been considered [81], but the approach is computationally expensive and consequently not practically feasible for the large-scale AF detection task considered here. Full details of the involved steps and their linkage, and how aleatoric and epistemic uncertainty are captured, can be found in [75].